Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labour-hours, and its standard costs per unit are as follows: Direct materials: 4 kg at $9.00 per kg $ 36.00 Direct labour: 3 hours at $12 per hour 36.00 Variable overhead: 3 hours at $8 per hour 24.00 Total standard cost per unit $ 96.00 The company planned to produce and sell 28,000 units in March. However, during March the company actually produced and sold 33,000 units and incurred the following costs: Purchased 165,000 kg of raw materials at a cost of $7.20 per kg. All of this material was used in production. Direct labour: 58,000 hours at a rate of $13 per hour. Total variable manufacturing overhead for the month was $729,060.

1. What is the materials price variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.).)

2. What is the materials quantity variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.).)

5. What is the labour rate variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Do not round intermediate calculations.)

6. What is the labour efficiency variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Do not round intermediate calculations.)

7. What is the variable overhead spending variance for March? (Do not round intermediate calculations. Round the actual overhead rate to two decimal places. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.).)

8. What is the variable overhead rate variance for March? (Round the actual overhead rate to two decimal places. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.).)
Answer #1

Answers

1.materials price variance for March

Material price Varience = (Standard price-Actual Price) * Actual Quantity

Actual Price = 7.2

Actual Quantity=165000

Standard Price=9

Material Price variance=(9 - 7.2) * 165000 =297000 F

2. materials quantity variance for March

Material Quantity Varience = (Standard Quantity -Actual Quantity) * Standard Price

Standard Quantity= 4 * 33000 = 132000

Actual Quantity= 165000

Standard Price=9

Material Quantity Varience = (132000 - 165000) * 9 = 297000 U

5. labour rate variance for March

Direct Labour rate variance = (Standard Rate - Actual Rate) * Actual Hours

Standard Rate = 12

Actual Rate = 13

Actual Hours = 58000

Direct Labour rate variance = (12 - 13) * 58000 = 58000 U

6. labour efficiency variance for March

Labour Efficiency Varience = (standard hour - Actual hour) * Standard rate

standard hour = 33000*3 = 99000

Actual hour = 58000

Standard rate = 12

Labour Efficiency Varience = (99000 - 58000) * 12 = 492000

7.variable overhead spending variance for March

variable overhead spending variance for = actual hours * (Actual manufacturing variable overhead rate - standard variable overhead rate )

actual hours = 58000

actual variable manufacturing overhead for the month was = 729060

standard Variable overhead = 8 per hour

actual variable overhead =  729060/58000 = 12.5

variable overhead spending variance for = 58000 (12.5 - 8) = 57992

8. variable overhead rate variance for March

variable overhead rate variance = (Standard rate - Actual rate)actual hours

Standard rate = 8

Actual rate = 729060/87000 = 8.38

Actual hours = 87000

variable overhead rate variance = (8 - 838) * 87000 = 33060 U